英語 (English)

基盤英語 (Basic English) (薬 (薬 B1 年))

早内-プリングルジュディス · 非常勤講師 1 単位前期火 5 ・ 6

(平成 19 年度以前の授業科目: 『基盤英語』 ) (平成 16 年度以前 (医保は 17 年度以前) の授業科目: 『英語 (1)』 )

【授業の目的】To give students opportunities to develop their English abilities through the exploration and discussion of a variety of themes. Students will develop their real-life communicative skills while expanding their vocabulary, grammar ability and fluency. A textbook will be used along with other items. Practice in the four skills (reading, writing, listening and speaking) will be provided during class and through homework.

【授業の概要】Students are expected to speak and write only in English during class. There will be homework/projects outside of class, and students are expected to complete those assignments even if they are absent. There will be presentations, etc. The following class schedule is tentative and may be modified due to factors such as class size, etc.

【キーワード】現代英語，コミュニケーション，外国文化【先行科目】 [先行科目] 【関連科目】 [関連科目] 【到達目標】In today’s globalized society, English has become the common

language of communication and is necessary to gain more knowledge of world events, and process information. This class requires students to put their English skills into practice, and they will increase in fluency and confidence as they do so.

【授業の計画】 1. General Class Guidance

2. Names

3. Kiss, Bow, or Shake Hands?

4. Prized Possessions

5. Cheers!

6. What’s the Occasion?

7. First Dates

8. Mid-term evaluation and presentation

9. On the Job

10. A Gift for Me?

11. Feast on This

12. Looking Good

13. That’s Shocking!

14. Glued to the Tube

15. Final Evaluation & Presentation 1

16. Final Evaluation & Presentation 2 【教科書】Linda Lee, Kensaku Yoshida & Steve Ziolkowski: J-Talk, Oxford

University Press 【参考書等】 [参考資料] 【成績評価の方法】 Weekly class participation, attendance, homework: 70%; Mid-term

& Final Evaluations: 30% 【再試験の有無】 Allowable if the student has attended more than 2/3 of the classes

and has a score of 50% or higher. 【受講者へのメッセージ】 [注意]

【授業コンテンツ】 http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=220839 【連絡先 (オフィスアワー・研究室・ E メールアドレス)】

⇒ 早内-プリングル (juditheph@yahoo.com) ���� (オフィスアワー: By appointment)

English

Basic English (薬 (薬 B1 年))

Judith Hayauchi · Part-time Lecturer 1 unit 前期火 5 ・ 6

(平成 19 年度以前の授業科目: 『基盤英語』 ) (平成 16 年度以前 (医保は 17 年度以前) の授業科目: 『英語 (1)』 )

Target⟩ To give students opportunities to develop their English abilities through the exploration and discussion of a variety of themes. Students will develop their real-life communicative skills while expanding their vocabulary, grammar ability and fluency. A textbook will be used along with other items. Practice in the four skills (reading, writing, listening and speaking) will be provided during class and through homework.

Outline⟩ Students are expected to speak and write only in English during class. There will be homework/projects outside of class, and students are expected to complete those assignments even if they are absent. There will be presentations, etc. The following class schedule is tentative and may be modified due to factors such as class size, etc.

�������⟩ present-day English, communication, foreign culture Fundamental Lecture⟩ [先行科目] Relational Lecture⟩ [関連科目] Goal⟩ In today’s globalized society, English has become the common language of

communication and is necessary to gain more knowledge of world events, and process information. This class requires students to put their English skills into practice, and they will increase in fluency and confidence as they do so.

Schedule⟩

1. General Class Guidance

2. Names

3. Kiss, Bow, or Shake Hands?

4. Prized Possessions

5. Cheers!

6. What’s the Occasion?

7. First Dates

8. Mid-term evaluation and presentation

9. On the Job

10. A Gift for Me?

11. Feast on This 12. Looking Good

13. That’s Shocking!

14. Glued to the Tube

15. Final Evaluation & Presentation 1

16. Final Evaluation & Presentation 2 Textbook⟩ Linda Lee, Kensaku Yoshida & Steve Ziolkowski: J-Talk, Oxford

University Press Reference⟩ [参考資料] Evaluation Criteria⟩ Weekly class participation, attendance, homework: 70%;

Mid-term & Final Evaluations: 30% Re-evaluation⟩ Allowable if the student has attended more than 2/3 of the classes

and has a score of 50% or higher. Message⟩ [注意]

Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=220839 Contact (Office-Hour, Room, E-mail)⟩

⇒ Hayauchi (juditheph@yahoo.com) ���� (Office Hour: By appointment)